OpenStudy (anonymous):

find the intercepts y=-2-2x

OpenStudy (anonymous):

lug x =0 solve for y lug y =0 solve for x

OpenStudy (anonymous):

y=-2-2(0) y=-2

OpenStudy (anonymous):

0=-2-2x 2x=-2 x=-2/2 x = -1

OpenStudy (anonymous):

If I deserve medal then give me one